Adam, Brenda, and Carl are running a relay race. Carl runs 1 yard per second slower
than Adam and Brenda runs 10% faster than Adam. They each run for 10 seconds and they
run a total of 145 yards. How fast does each run?

Found 2 solutions by ikleyn, MathTherapy:
Answer by ikleyn(52835)

.
Adam, Brenda, and Carl are running a relay race. Carl runs 1 yard per second slower
than Adam and Brenda runs 10% faster than Adam. They each run for 10 seconds and they
run a total of 145 yards. How fast does each run?
~~~~~~~~~~~~~~~~~~

Let A = rate of Adam, in yards per second.

Then the Carl's rate is (A-1) yards per second, while Brenda's rate is 1.1A yerds per second.

In 10 seconds Adam   runs 10A yards,
              Brenda runs 10*1.1A yards,
              Carl   runs 10*(A-1) yards.

In total,

10A + 10*1.1A + 10*(A-1) = 145.

Simplify:

10A + 11A + 10A - 10 = 145,

31A = 145 + 10,

31A = 155.


Solve for A and then complete the solution on your own.


Answer by MathTherapy(10555)

Adam, Brenda, and Carl are running a relay race. Carl runs 1 yard per second slower
than Adam and Brenda runs 10% faster than Adam. They each run for 10 seconds and they
run a total of 145 yards. How fast does each run?
Let Adam’s speed be S
Then Carl’s speed is: S – 1, and Brenda’s is: 1.1S
We then get the following DISTANCE equation: 10(S) + 10(S – 1) + 10(1.1S) = 145
10S + 10S – 10 + 11S = 145
31S = 145 + 10
31S = 155
S, or